# A model-independent measurement of the expansion and growth rates from BOSS using the FreePower method

### v2

@@ 1 Introduction
An approach to remain model independent, which was recently proposed in [27, 29, 30], consists in dividing the linear matter power spectrum $P(k)$ into several $k$-bins whose values are free to vary. From these bins, and a set of bias and counterterm parameters, the non-linear galaxy power spectrum and bispectrum are derived, allowing comparisons with the galaxy survey data. In this case, it is not necessary to make assumptions about the shape of $P(k)$. The non-linear correction to one loop in the power spectrum are evaluated assuming very general kernels, derived under the assumption of a homogeneous and isotropic background and the equivalence principle of general relativity, which implies that one can remove a pure-gradient metric perturbation by going to the free-falling frame of comoving observers [24]. All the parameters are left free to vary in each redshift. In this way, the results are independent of the details of both early-time models (that determine the initial power spectrum shape) and late-time models (that determine the background and perturbations growth). In Ref. [30] this framework was denoted as the FreePower method.
@@
The main aim of this work is to produce the first real-data analysis with the FreePower framework using data from the Baryon Oscillation Spectroscopic Survey (BOSS). We do not strive for the best possible precision in this first real-data application of FreePower. Instead we focus on demonstrating that the method works and is able to produce competitive constraints for the $E(z)$ and $f(z)$. Since each likelihood evaluation is expensive and, as discussed below, we will have 17 free parameters, we had difficulty achieving convergence of the MCMC chains when using a small, 256-core computing cluster. We thus adopted three main simplifications in order to reduce the total computational time and get reliable results in a reasonable amount of time. First, we do not include the bispectrum, which if included would result in increased precision [30]. Second, we adopt the analytical kernels for an Einstein-deSitter (EdS) model. EdS kernels should in any case be a very good approximation for cosmologies that do not depart too much from $\Lambda$CDM; they have been adopted also in a previous analysis of BOSS data that we will compare to [14]. Finally, we include an informative distance prior, which we chose as Gaussians with $3\%$ standard deviations centered on the best-fit Planck $\Lambda$CDM values for each redshift bin.
@@ 2 The FreePower model-independent approach
The theoretical model for the the galaxy density field power spectrum is given as [14, 35]
@@
This spectrum includes several free bias parameters [36], denoted $b_{1}$, $b_{2}$, $b_{\mathcal{G}_{2}}$, and $b_{\Gamma_{3}}$, and two counterterm parameters $c_{0},c_{2}$. In Ref. [30] we also included an overall smoothing factor that takes into account the Finger-of-God effect and the redshift uncertainty; we found however that the effect of this smoothing was subdominant and here, therefore, for simplicity, we do not include it. More details can be found in Appendix C. We explore the 17-dimensional parameter space with the Python MCMC code emcee [37]. [^1] https://emcee.readthedocs.io/en/stable/ (last access: 21.06.2024)
In the model-independent approach presented here the linear matter power spectrum is parametrized in $k$-bins and the waveband values are varied in the MCMC. We select the values of $P_{L}$ at the central $k$-values
@@
Previous tests indicate that neither the exact spacing nor the number of the $k$-bins has a strong influence on the results on the other parameters [38]. The linear matter power spectrum is constructed with an interpolation between these varied parameters. Because of that, the linear matter power spectrum is quite smooth. Therefore, the coupling between the large-scale IR modes and the mildly non-linear modes is rather weak [10] and the effect of IR resummation is in general small. In this approach we do not disentangle the growth factor $G(z)$ from the linear power spectrum, so we in fact measure $P_{L}(k,z)=G(z)^{2}P_{L}(k)$ at each redshift bin.
In the present analysis, we do not use the hexadecapole because it has been shown that the signal-to-noise ratio in the hexadecapole is very low compared to monopole and quadrupole [36]. [^2] Hexadecupole functionality was also absent from PyBird until very recently, already in the late stages of our analysis. Therefore, the model prediction for the hexadecapole is set to zero. However, due to the fact that at most monopole and quadrupole are considered in the likelihood in the final analysis, this simplification can only influence the possible coupling between multipoles through the wide-angle effects [39] and this effect should be small. For the data comparison, the $k$-range is restricted to $0.01\,h/{\rm Mpc}<k<0.3\,h/{\rm Mpc}$.
The bias parameters $b_{1}$, $b_{2}$, and $b_{\mathcal{G}_{2}}$ are left free to vary, together with the counterterm parameters $c_{0}$ and $c_{2}$. The bias parameter $b_{\Gamma_{3}}$ is set to zero, following [14], since it was observed that it is degenerate with the counterterm parameters and $b_{\mathcal{G}_{2}}$. The parameter $c_{4}$ need not be considered since we do not use the hexadecapole here. The parameters are in the so-called EC basis [36]. We also vary the growth rate $f$ and the dimensionless shot noise parameter
$s=P_{\rm shot}\bar{n}-1\,,$ (2.3)
where $\bar{n}\approx 3\times 10^{-4}(h/{\rm Mpc})^{3}$ [14] is the number density of the galaxies in BOSS.
@@ 2.1 Likelihood
$\mathcal{L}\propto\exp[-\frac{1}{2}(\mathbf{P}_{\rm dat}-WM\mathbf{P}_{\rm theo})^{T}C^{-1}(\mathbf{P}_{\rm dat}-WM\mathbf{P}_{\rm theo})\bigg{]}\,.$ (2.4)
$\mathbf{P}_{\rm dat}$ contains the data for the power spectrum multipoles $l=0,1,2,3,4$ in the $k$-range $0-0.4\,h/{\rm Mpc}$ in bins of $0.01\,h/{\rm Mpc}$. It has a vector of length $200$, combining these five multipoles and 40 $k$ bins. However, it is important to note that any analysis should be limited to $k<0.3\,h/{\rm Mpc}$ due to the Nyquist-frequency in the measured power spectra [39] and because the theoretical one-loop spectrum is hardly reliable at smaller scales. Therefore, we use $k_{\rm max}=0.3\,h/{\rm Mpc}$. $\mathbf{P}_{\rm theo}$ is a vector that contains the even multipoles of the model in $k$-bins of $0.001\,h/{\rm Mpc}$. $W$ is a 200$\times$2000 survey window function matrix and $M$ is the 1200$\times$2000 wide-angle transformation matrix [39]. We compare the theoretical predictions of the model to monopole and quadrupole of the data vector.
@@ 2.2 Data sets from BOSS
linear growth rate $f$ ${\cal U}[0,1]$ ${\cal U}[0,1.5]$
@@
dim.less angular diameter distance $L_{A}$ $\mathcal{N}(0.25,0.008)$ $\mathcal{N}(0.322,0.010)$
@@
The effective redshift can be used because the redshift bin is much smaller than the scale of variation of the growth factor $G(z)$ and the non-cosmological parameters. Therefore, the binning corresponds to evaluating all quantities at the same effective redshift [15]. Due to the time constraints on our available computing system, in this study only the (larger) NGC data set at both redshift bins is employed. More specifically, we compare our one-loop spectrum to the BOSS NGC monopole and quadrupole spectra.
@@
The priors of the varied parameters are presented in Table 1. The priors on the eight parameters describing the linear matter power spectrum, as well as the bias parameters and counterterms were originally intended to be very broad and non-informative. However, the slow convergence of the chains prompted us to use tighter priors. For $P_{i}$ we decided to set them by running a MCMC with 256 walkers for 10000 steps for the high-$z$ case with flat $P$ priors and then doubling the standard deviation of the obtained posteriors for $P_{3}$-$P_{8}$ and adopting a tighter $10\%$ prior for $P_{1}$ and $P_{2}$ since those bins are on scales larger than the survey window. This is intended as a balance between using loose priors and helping MCMC convergence. As is shown in Appendix B, the $P_{3}$-$P_{8}$ priors remain considerably broader than the posteriors. For the shot-noise parameter, we adopt Gaussian priors which allows for a conservative 50% effective variation of the galaxy number density $\bar{n}$.
@@
For the bias parameters, we employed Gaussian priors using as reference the final NGC posterior results from ISZ20. We took their errorbars, propagated the uncertainty in the power spectrum amplitude $A$ since they multiply all their bias parameters by $(A/A_{\rm Planck})^{1/2}$, symmetrized the results and multiplied them by a factor of four. We used the final results for the mean and standard deviation to construct our low-$z$ uncorrelated Gaussian priors. We also did a similar procedure for the counterterm parameters, but only multiplied the results by a factor of two. We thus allow a four times larger uncertainty of ISZ20 in each bias and twice as large for the counterterm parameters as our prior. This resulted in posteriors which were never driven by the priors themselves. I.e., these priors are weakly informative. The Gaussian prior on $L_{A}(z)$ instead is informative and is centered around its Planck 2018 $\Lambda$CDM value (i.e. assuming $\Omega_{m0}=0.31$ so that $L_{A}=0.25$ for the low-$z$ and 0.322 for the high-$z$ sample) with a $3\%$ standard deviation. For $z\leq 1$ current constraints from Pantheon+ [46] are already at this level.
@@ 3 MCMC results
Figure 1: $68.3\%$ and $95.4\%$ highest-density confidence intervals corner plot for the cosmological and bias terms for the red low-$z$ sample ($z_{\rm eff}=0.38$). Orange dots mark the Planck 2018 $\Lambda$CDM values. The orange curve represents the $L_{A}$ vs. $E$ relation for $\Lambda$CDM for different values of $\Omega_{m0}$. $L_{A}$ results are prior-dominated.
@@
We ran emcee using 256 walkers initialized in a small volume in parameter space. We ran, for each walker, 800k steps, for a total of 200 million chain points for each low-$z$ and high-$z$ samples. Due to the complexity of the 1-loop calculations and the amount of MCMC steps needed, the total computational time used for both redshift bin runs was large, around 100k CPU-hours. This heavy computation cost was the main reason for our choice of a small number of $k$-bins, limiting somewhat the dimensionality of our parameter space. We analyzed the results using both the Python package getdist [47] and our own codes. We compared our final results with those obtained with smaller chains with 400k steps to confirm the results we found were robust, which as we will discuss is in tension with $\Lambda$CDM CMB results. Indeed, the results with half the chain steps change little, which gives us further confidence in the convergence.
@@
From the $P(k)$ linear bins we can also obtain the amplitude $\sigma_{8}$ by interpolating and integrating over a spherical window function. We find $\sigma_{8}$ to be $0.89\pm 0.08$ ($0.82\pm 0.08$) for the $z_{\rm eff}=0.38$ ($0.61$) redshift bin. Note, however, that due to the small number of $k$-bins here considered, the value of $\sigma_{8}$ will inevitably depend on the chosen interpolation scheme. For a standard $\Lambda$CDM-like $P(k)$ binned in the same way as our spectra, our tests show that our chosen interpolation can recover the fiducial $\sigma_{8}$ to within 0.02. Our results are around $1\sigma$ higher than Planck 2018. They also change by less than 2% when discarding the first two prior-dominated $k$ bins.
@@
The correlations $\rho$ among the 17 parameters are in most cases small. The average off-diagonal absolute value of $\rho$ is around 0.16 for both redshift bins. We list here some particular values for $z_{\rm eff}=0.38$ (0.61): $\rho(f,E)=0.59$ ($0.73$), $\rho(E,L_{A})=-0.31$ ($-0.31$), $\rho(E,b_{1})=-0.32$ ($-0.57$), $\rho(c_{0},c_{2})=0.75$ ($0.44$). There is some large correlations between individual $k$-bins of $P(k)$ and some bias parameters. But both $f$ or $E$ show only small or moderate correlations with $P(k)$ and the bias parameters. The full correlation matrices for both bins are shown in Appendix A.
@@
We tested the convergence of our MCMC using primarly the auto-correlation time ($\tau$) estimates [37]. A large ratio between the number of steps and $\tau$ indicate that the chain ran for long enough to have a large effective number of independent steps. Ideally $\tau$ should converge to an asymptotic value for high enough number of steps. We found that around 400k steps with each of 256 walkers were needed for reaching convergence. More details are provided in Appendix A. We also performed the Gelman-Rubin test [49]. This simple test is not the most reliable, specially using an ensemble sampler like emcee since in this approach the walkers are not independent. Nevertheless it gives another handle on convergence which is independent of the auto-correlation time analysis. For the cosmological parameters $f$, $E$ and $L_{A}$, we get $(\hat{R})^{(1/2)}=\{1.028,1.044,1.016\}$ for the low-$z$ run, and $(\hat{R})^{(1/2)}=\{1.035,1.042,1.018\}$ for the high-$z$ run. It is commonly assumed as a threshold for convergence the value $(\hat{R}_{\rm max})^{(1/2)}=1.1$. This agrees qualitatively with the auto-correlation time results in Appendix A, which indicate reasonable convergence for the chains.

### v3

@@ 1 Introduction
An approach to remain model independent, which was recently proposed in [27, 29, 30], consists in dividing the linear matter power spectrum $P(k)$ into several $k$-bins whose values are free to vary. From these bins, and a set of bias and counterterm parameters, the non-linear galaxy power spectrum and bispectrum are derived, allowing comparisons with the galaxy survey data. In this case, it is not necessary to make assumptions about the shape of $P(k)$. At the same time, the expansion rate $H(z)$ and the growth rate $f(z)$ are left free to vary with redshift. The non-linear correction to one loop in the power spectrum are evaluated assuming very general kernels, derived under the assumption of a homogeneous and isotropic background and the equivalence principle of general relativity, which implies that one can remove a pure-gradient metric perturbation by going to the free-falling frame of comoving observers [24]. All the non-linear parameters are left free to vary in each redshift. In this way, the results are independent of the details of both early-time models (that determine the initial power spectrum shape) and late-time models (that determine the background and perturbations growth). In Ref. [30] this framework was denoted as the FreePower method.
@@
The main aim of this work is to produce the first real-data analysis with the FreePower framework using data from the Baryon Oscillation Spectroscopic Survey (BOSS). We do not strive for the best possible precision in this first real-data application of FreePower. Instead we focus on demonstrating that the method works and is able to produce competitive constraints for the $E(z)$ and $f(z)$. Since each likelihood evaluation is expensive and, as discussed below, we will have 17 free parameters, we had difficulties with the large computational cost of the MCMC codes when using a small, 256-core computing cluster. We thus adopted two main simplifications with respect to [30] in order to reduce the total computational time and get reliable results in a reasonable amount of time. First, we do not include the bispectrum, which if included would result in increased precision [30]. Second, we adopt the analytical kernels for an Einstein-deSitter (EdS) model. EdS kernels should in any case be a very good approximation for cosmologies that do not depart too much from $\Lambda$CDM; they have been adopted also in a previous analysis of BOSS data that we will compare to [14].
@@ 2 The FreePower model-independent approach
2.1 Theoretical model
The theoretical model for the galaxy density field power spectrum is given as [14, 35]
@@ 2.1 Theoretical model
The summands are the tree-level galaxy power spectrum, the two one-loop corrections, the countertems, and the shot noise. This spectrum includes model parameters describing the galaxy bias [36], counterterms, and the shot noise through the dimensionless shot noise parameter
$s=P_{\rm shot}\bar{n}-1\,,$ (2.2)
where $\bar{n}\approx 3\times 10^{-4}(h/{\rm Mpc})^{3}$ [14] is the number density of the galaxies in BOSS. The parameters are in the so-called EC basis [36]. The model also depends on the cosmological parameters $E$, $L_{A}$, and the growth rate $f$.
In Ref. [30] we also included an overall smoothing factor that takes into account the Finger-of-God effect and the redshift uncertainty; we found however that the effect of this smoothing was subdominant and here, therefore, for simplicity, we do not include it. More details regarding the individual terms can be found in Appendix C. Importantly, the model (2.1) uses the linear matter power spectrum which, in standard analyses, is calculated within specific cosmological models. In contrast, in the model-independent approach presented here, the linear matter power spectrum is parametrized in $k$-bins and the waveband values are varied in the MCMC. We select the values of $P_{L}$ at the central $k$-values
@@
In this way, we introduce eight new parameters to vary in the MCMC analysis but there is no underlying assumption about the cosmological model when the linear matter power spectrum is calculated anymore. Previous tests indicate that neither the exact spacing nor the number of the $k$-bins has a strong influence on the results on the other parameters [37]. The linear matter power spectrum is constructed with an interpolation between these varied parameters. Because of that, the linear matter power spectrum is quite smooth. Therefore, the coupling between the large-scale IR modes and the mildly non-linear modes is rather weak [10] and the effect of IR resummation is in general small. In this approach we do not disentangle the growth factor $G(z)$ from the linear power spectrum, so we in fact measure $P_{L}(k,z)=G(z)^{2}P_{L}(k)$ at each redshift bin.
In this application of the FreePower method, the theoretical power spectrum (2.1) depends on a number of free parameters at each redshift (see Table 1): the eight linear power spectrum wavebands (2.3); the background functions $H(z),L_{A}(z)$ that enter the Alcock-Paczynski effect; the linear growth rate $f(z)$; the linear and non-linear bias parameters $b_{1},b_{2},b_{\mathcal{G}_{2}}$; the counterterms $c_{0},c_{2}$; and the shot noise $s$ (see App. C for the definition of these parameters). They are not parametrized in terms of a cosmological model like e.g. $\Lambda$CDM but taken as fully free fundamental parameters in the Fisher matrix, as we explain next.
@@ 2.2 Likelihood
$\mathcal{L}\propto\exp[-\frac{1}{2}\sum_{\ell={0,2}}\sum_{\ell^{\prime}={0,2}}\Big{(}\mathbf{P}_{\rm dat,\ell}-(WM\mathbf{P}_{\rm theo})_{\ell}\Big{)}^{T}C^{-1}_{\ell\ell^{\prime}}\Big{(}\mathbf{P}_{\rm dat,\ell^{\prime}}-(WM\mathbf{P}_{\rm theo})_{\ell^{\prime}}\Big{)}\Bigg{]}\,.$ (2.4)
The full BOSS dataset contains five different multipoles: $\ell=0,1,2,3,4$ in the $k$-range $0-0.4\,h/{\rm Mpc}$. It has 40 $k$ bins and a total vector of length $200$ (40 for each of the five multipoles). $W$ is a 200$\times$2000 survey window function matrix and $M$ is the 1200$\times$2000 wide-angle transformation matrix which transforms 3 multipoles into the 5 final multipoles [38]. $\mathbf{P}_{\rm theo}$ is a vector that contains the even multipoles ($\ell=0,2,4$) of the model in $k$-bins of $0.001\,h/{\rm Mpc}$. For our analysis, we choose $k_{\rm max}=0.3\,h/{\rm Mpc}$ as our default limit. This is motivated by large volume simulations which have shown that one-loop is still reliable up to this scale for BOSS data [14, 39].111 We remark that any analysis should be limited to $k<0.3\,h/{\rm Mpc}$ due to the Nyquist-frequency in the measured power spectra [38]. We also neglect $\ell=1,3,4$. So our $\mathbf{P}_{\rm dat}$ contains the data for the power spectrum multipoles $\ell=0,2$ in the $k$-range $0-0.3\,h/{\rm Mpc}$ in bins of $0.01\,h/{\rm Mpc}$.
@@ 2.3 Data sets from BOSS
linear growth rate $f$ ${\cal U}[0,1.5]$ ${\cal U}[0,1.5]$
@@
dim.less angular diameter distance $L_{A}$ ${\cal U}[0.12,0.38]$ ${\cal U}[0.16,0.48]$
@@
The effective redshift can be used because the redshift bin is much smaller than the scale of variation of the growth factor $G(z)$ and the non-cosmological parameters. Therefore, the binning corresponds to evaluating all quantities at the same effective redshift [15].
2.4 Settings for our analysis
Due to the time constraints on our available computing system, in this study only the (larger) NGC data set at both redshift bins is employed.
We compare the theoretical predictions of the one-loop power spectrum model to monopole and quadrupole of the data vector. We only calculate the $2\times 400$ values of monopole and quadrupole of the model vector $\mathbf{P}_{\rm theo}$ and only use the $2\times 40$ values of monopole and quadrupole of $WM\mathbf{P}_{\rm theo}$ to compare them to the data vector $\mathbf{P}_{\rm dat}$ in the likelihood. The odd multipoles and the hexadecapole in $\mathbf{P}_{\rm dat}$ are not used.
It has been shown that the signal-to-noise ratio in the hexadecapole is very low compared to monopole and quadrupole [36]. [^3] Hexadecupole functionality was also absent from PyBird until very recently, already in the late stages of our analysis. Therefore, the model prediction for the hexadecapole is set to zero. However, due to the fact that at most monopole and quadrupole are considered in the likelihood, this simplification can only influence the possible coupling between multipoles through the wide-angle effects [38] and this effect should be small. For the data comparison, the $k$-range is restricted to $0.01\,h/{\rm Mpc}<k<0.3\,h/{\rm Mpc}$.
The bias parameters $b_{1}$, $b_{2}$, and $b_{\mathcal{G}_{2}}$ are left free to vary, together with the counterterm parameters $c_{0}$ and $c_{2}$. The bias parameter $b_{\Gamma_{3}}$ is set to zero, following [14], since it was observed that it is degenerate with the counterterm parameters and $b_{\mathcal{G}_{2}}$. The parameter $c_{4}$ need not be considered since we do not use the hexadecapole here.
@@ 2.4 Settings for our analysis
The priors of the varied parameters are presented in Table 1. The priors on the eight parameters describing the linear matter power spectrum, as well as the bias parameters and counterterms were originally intended to be very broad and non-informative. However, the slow convergence of the chains prompted us to use tighter priors. For $P_{i}$ we decided to set them by using a previously run small MCMC chain using emcee with 256 walkers for 10000 steps for the high-$z$ case with flat $P$ priors and then doubling the standard deviation of the obtained posteriors for $P_{3}$-$P_{8}$ and adopting a tighter $10\%$ prior for $P_{1}$ and $P_{2}$ since those bins are on scales larger than the survey window. This is intended as a balance between using loose priors and helping MCMC convergence. As is shown in Appendix B, the $P_{3}$-$P_{8}$ priors remain reasonably broader than the posteriors. For the shot-noise parameter, we adopt Gaussian priors which allows for a conservative 50% effective variation of the galaxy number density $\bar{n}$.
@@
For the bias parameters, we employed Gaussian priors using as reference the final NGC posterior results from ISZ20. We took their errorbars, propagated the uncertainty in the power spectrum amplitude $A$ since they multiply all their bias parameters by $(A/A_{\rm Planck})^{1/2}$, symmetrized the results and multiplied them by a factor of four. We used the final results for the mean and standard deviation to construct our low-$z$ uncorrelated Gaussian priors. We also did a similar procedure for the counterterm parameters, but only multiplied the results by a factor of two. We thus allow a four times larger uncertainty of ISZ20 in each bias and twice as large for the counterterm parameters as our prior. This resulted in posteriors which were never driven by the priors themselves. I.e., these priors are weakly informative.
For both $E$ and $L_{A}$ we take very large and uninformative uniform priors, centered around their respective $\Lambda$CDM value. We notice that current SN constraints on $L_{A}$ are already much tighter than this (for instance, taking the 16 SN in the redshift range $z\in(0.37,0.39)$ from the Pantheon+ [46] catalog produces an error around 3%). However, their use would need the assumption of standard candles, while in this paper we prefer to focus exclusively on clustering data.
@@ 3 MCMC results
We explore the 17-dimensional parameter space with the Python MCMC codes emcee [47] [^4] https://emcee.readthedocs.io/en/stable (last access: 21.06.2024) and ptemcee [48]. [^5] https://github.com/willvousden/ptemcee (last access: 21.11.2024) We found that emcee exhibited very large autocorrelation times ($\sim 10000$) for most parameters, which led to a very slow convergence time even when using 256 walkers. We therefore resorted to the modified version ptemcee, which uses parallel tempering, for which the autocorrelation time was typically only $\sim 10$. All of our following results were obtained using ptemcee.
Figure 1: $68.3\%$ and $95.4\%$ highest-density confidence intervals corner plot for the cosmological and bias terms for the red low-$z$ sample ($z_{\rm eff}=0.38$). Orange dots mark the Planck 2018 $\Lambda$CDM values. The orange curve represents the $L_{A}$ vs. $E$ relation for $\Lambda$CDM for different values of $\Omega_{m0}$.
@@
We ran ptemcee using 60 walkers and 6 different temperatures, initialized in a large volume in the allowed prior space. We ran, for each walker, 300k steps, for a total of 108 million chain points for each low-$z$ and high-$z$ samples. Due to the complexity of the 1-loop calculations and the amount of MCMC steps needed, the total computational time used for both redshift bin runs was large, around 50k CPU-hours. This heavy computation cost was the main reason for our choice of a small number of $k$-bins, limiting somewhat the dimensionality of our parameter space. We analyzed the results using both the Python package getdist [49] and our own codes. We compared our results obtained with both ptemcee and emcee to confirm the results we found were robust, which as we will discuss is in tension with $\Lambda$CDM CMB results.
@@
From the $P(k)$ linear bins we can also obtain the amplitude $\sigma_{8}$ by interpolating and integrating over a spherical window function. We find $\sigma_{8}$ to be $0.887\pm 0.094$ ($0.799\pm 0.083$) for the $z_{\rm eff}=0.38$ ($0.61$) redshift bin. Note, however, that due to the small number of $k$-bins here considered, the value of $\sigma_{8}$ will inevitably depend on the chosen interpolation scheme. For a standard $\Lambda$CDM-like $P(k)$ binned in the same way as our spectra, our tests show that our chosen interpolation can recover the fiducial $\sigma_{8}$ to within 0.02. Our results are around $1\sigma$ higher than Planck 2018. They also change by less than 2% when discarding the first two prior-dominated $k$ bins.
@@
The correlations $\rho$ among the 17 parameters are in most cases small. The average off-diagonal absolute value of $\rho$ is around 0.17 for both redshift bins. We list here some particular values for $z_{\rm eff}=0.38$ (0.61): $\rho(f,E)=-0.32$ ($-0.12$), $\rho(E,L_{A})=-0.80$ ($-0.69$), $\rho(E,b_{1})=-0.75$ ($-0.76$), $\rho(L_{A},b_{1})=0.83$ ($0.75$), $\rho(c_{0},c_{2})=0.65$ ($0.41$). There is some large correlations between individual $k$-bins of $P(k)$ and some bias parameters. But $f$, $E$ and $L_{A}$ show only small or moderate correlations with $P(k)$ and the bias parameters. The full correlation matrices for both bins are shown in Appendix A.
@@
We tested the convergence of our MCMC using primarly the auto-correlation time ($\tau$) estimates [47]. A large ratio between the number of steps and $\tau$ indicate that the chain ran for long enough to have a large effective number of independent steps. $\tau$ should converge to an asymptotic value for high enough number of steps. We found that using ptemcee $\sim 1000$ steps were necessary to reach convergence, while using emcee one needed almost a million steps. We also performed the simple Gelman-Rubin test [51]. Both results indicate complete convergence of the chains. More details are provided in Appendix A.